Opened 9 years ago

Closed 8 years ago

Last modified 10 months ago

#282 closed flight processing (fixed)

ET07/01, flight day 298/2009, Afar mid

Reported by: benj Owned by: knpa
Priority: alpha 5 Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors: knpa

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/ethiopia/ET07_01-2009_298_Afar_mid

Data arrived from ARSF via SATA disk Y on 4th Nov. 2009.

Scientific objective: /users/rsg/arsf/arsf_data/2008/ARSF_Applications/Ethiopia/ET07_01\ -\ Dr\ Wright.pdf

Priority: alpha-5

PI: Tim Wright

2009 Ethiopia campaign, reflown from 2007

Sensors:

  • Eagle (09/02/10)
  • Hawk (09/02/10)
  • Leica LIDAR (09/02/10)
  • RCD (09/02/10)
  • TABI (to be processed by University of Athens)

Change History (31)

comment:1 Changed 9 years ago by benj

Raw GPS base station data is in ~arsf/arsf_data/2009/flight_data/ethiopia/basestations.

comment:2 Changed 9 years ago by mggr

  • Description modified (diff)

Raw TABI data acquired, put in archive and offered to Athens for download and processing.

comment:3 Changed 9 years ago by mggr

LIDAR and webcam imagery disentangled. Unrelated LIDAR files removed.

comment:4 Changed 9 years ago by benj

  • Owner set to benj
  • Status changed from new to assigned

comment:5 Changed 9 years ago by benj

Eagle lines 4 and 9, hdr file recorded that there was only 1 line. Data itself appears OK, corrected number of lines in hdr file.

comment:6 Changed 9 years ago by knpa

LIDAR lines processed, QC underway.

comment:7 Changed 9 years ago by knpa

  • Other processors set to knpa

comment:8 Changed 9 years ago by knpa

LIDAR processing complete, preparing delivery.
Having trouble with DEM creation.

comment:9 Changed 9 years ago by benj

Timing corrections for Eagle/Hawk. Note these are approximate because I've had to do them by comparison to the Lidar intensity image rather than by comparison to straight features in the E/H images themselves (since there aren't any).

Line Eagle Hawk
1 -0.05 -0.05
2 -0.05 0.00
3 0.07 -0.04
4 0.00 -0.04
5 -0.05 0.05
6 -0.10 -0.02
7 0.00 -0.01
8 -0.01 -0.05
9 0.00 -0.10
10 -0.07 0.00
11 -0.03 -0.10

comment:10 Changed 9 years ago by benj

Hyperspectral delivery is prepared and ready for delivery checking at ~airborne/workspace/ET07_01-2009_298_Afar_mid/delivery/20100202/ET07_01/

comment:11 Changed 9 years ago by mark1

Photographs 25025523100002G4.tif-25025523100006G4.tif have no time info in the logfile - times have been estimated from the sensorstat log to give approx nav/pos info for tagging.

comment:12 Changed 9 years ago by benj

Doing Lidar delivery check

comment:13 Changed 9 years ago by mark1

Photographs 25064951100001G4.tif to 25064951100069G4.tif have no time stamps in log file so processing them as if there was no logfile.

comment:14 Changed 9 years ago by benj

Lidar delivery looks fine, can be rsynced back and delivered.

comment:15 Changed 9 years ago by knpa

E+H delivery checked.

All bands look good apart from Hawk 233 (blank for all lines)
Unix2dos'd Readme.
Need to mention Eagle dropped scan lines in Readme.
Todo at some point: Readme should really clarify why there are sometimes HDF and Bil and the difference.
Todo at some point: Decided what to do about DEM - include a Readme to say may not be accurate?

comment:16 Changed 9 years ago by benj

Problem with last Hawk line being blank is a known issue (see #291).

Made remaining changes to readme.

comment:17 Changed 9 years ago by knpa

Repository full so skipping rsync for now, All sensors will need to be rsynced at a later date (will keep ticket open)

comment:18 Changed 9 years ago by knpa

All three projects copied on to external hard disk. Noticed some problems with the hyperspectral screenshots which need to be fixed before disk can be sent out.

comment:19 Changed 9 years ago by knpa

HYPERSPECTRAL, LIDAR AND PHOTOS DELIVERED TODAY (09/02/10)
Leaving this ticket open until rsynced.

comment:20 Changed 9 years ago by benj

Processed TABI data was received from Pol back in January but we weren't happy with the accuracy. This data is in the workspace, it has not been dispatched to the PI but is the only processed TABI data we currently have for this project.

comment:21 Changed 9 years ago by knpa

  • Description modified (diff)

comment:22 Changed 9 years ago by knpa

r-synced lidar

comment:23 Changed 9 years ago by knpa

  • Description modified (diff)

RCD data too!

comment:24 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from benj to knpa

Started archiving.

comment:25 Changed 9 years ago by knpa

Tarball complete and made available to NEODC.

comment:26 Changed 9 years ago by knpa

Downloaded and backed up by NEODC.

Moved to archive disk.

comment:27 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:28 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

comment:29 Changed 10 months ago by wja

New lidar delivery created with the addition of LAS files. ASCII and LAS files are now stored in 'flightlines' directory.

comment:30 Changed 10 months ago by wja

Updated table in Read_Me to include updated 'flightlines' directory structure along with inclusion of LAS files.

comment:31 Changed 10 months ago by wja

LAS path name (flightlines/las1.0/) changed to correct LAS format version (flightlines/las1.2). Read Me updated accordingly.

Note: See TracTickets for help on using tickets.